OpenStudy (anonymous):

giving medal for answer What is the total number of diagonals in an octagon? a. 9 c. 20 b. 5 d. 3

OpenStudy (anonymous):

i thought 20.?

OpenStudy (anonymous):

20 diagonals...... i have a formula... Number of diagonals= [n(n-3)] /2 n=sides octagon has 8 sides its equals to [8(8-3)] /2 = [8(5)] /2 = 40/2 = 20

OpenStudy (anonymous):

ok thnks just makin sure

OpenStudy (anonymous):

ok